Spider-Man is webslinging after the Green Goblin when the villain suddenly makes a sharp U turn to avoid capture. Having only a fraction of a second to react, Spider-Man fires a webline toward a telephone pole. Once the webbing is anchored, the webline enables Spider-Man to make a 3.85 m diameter U turn in 1.99 seconds. If Spider-Man's mass is 73.2 kg and the webline is the only force supplying the necessary centripetal force, what must be the tension in the webline (in newtons)?
Answer #1

To cover a u turn of diameter the time taken = 1.99 s

The distance covered = π r

r = radius = 3.85/2 = 1.925 m

Time period =2× 1.99 s

Angular speed w = π/1.99 rad/s

v = wr= 1.925π/1.99 = 3.039 m/s

Tension = mv^2 / r

T = 73.2*3. 039^2/1.925 = 351.19 N